Can you help me solve this problem: Using Microsoft Excel to calculate and analyze the Net Present Value (NPV), Internal Rate of Return (IRR), Accounting Rate of Return and Payback Period of the following investment.

Initial Investment $12,950
Estimated Life 10 years
Annual Cash Inflows $3,000
Cost of Capital 12%
Answer #1

NPV and IRR are calculated using NPV and IRR functions in Excel

NPV is $4,001

IRR is 19%

Accounting rate of return = annual cash inflow / initial investment = $3000 / $12950 = 23%

Payback period is the time taken for the cumulative cash flows to equal zero

Payback period = 4 + (cash flow required in year 5 for cumulative cash flows to equal zero / year 5 cash flow) = 4 + ($950 / $3000) = 4.32 years